OpenStudy (anonymous):

The system x+2y=5 3y+z=1 2x-z=8 can be represented by a matrix equation. What will be the dimensions of the coefficient matrix? Explain. Thank you:)

OpenStudy (anonymous):

\[\left[\begin{matrix}1 & 2 & 0\\ 0 & 3 & 1\\ 2 & 0 & -1\end{matrix}\right]\ \left[\begin{matrix}x \\ y \\z\end{matrix}\right] = \left[\begin{matrix}5 \\ 1 \\ 8\end{matrix}\right]\]

OpenStudy (anonymous):

not sure which parts needs explanations so please ask :)

OpenStudy (anonymous):

Thank you for showing me the matrix. I needed to find out the matrix dimension that was formed from that system of equations, and now I am able to answer that question. Thank you so veeeerrry much! =)

OpenStudy (anonymous):

glad i could help ^_^ don't hesitate to ask follow-up questions

OpenStudy (anonymous):

Wow! Thanks you're awesome!! =)